### v1

# ThinK: Thinner Key Cache by Query-Driven Pruning

Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications by leveraging increased model sizes and sequence lengths. However, the associated rise in computational and memory costs poses significant challenges, particularly in managing long sequences due to the quadratic complexity of the transformer attention mechanism. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence lengths, we uncover that the channel dimension of the KV cache exhibits significant redundancy, characterized by unbalanced magnitude distribution and low-rank structure in attention weights. Based on these observations, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in memory costs by over 20% compared with vanilla KV cache eviction methods. Extensive evaluations on the LLaMA3 and Mistral models across various long-sequence datasets confirm the efficacy of ThinK, setting a new precedent for efficient LLM deployment without compromising performance. We also outline the potential of extending our method to value cache pruning, demonstrating ThinK’s versatility and broad applicability in reducing both memory and computational overheads.
1 Introduction
Large language models (LLMs) (Hadi et al.,, 2023; Brown et al.,, 2020; OpenAI,, 2023; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Scao et al.,, 2022; Reid et al.,, 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks. A key principle, the Scaling Law (Kaplan et al.,, 2020), suggests that LLMs exhibit emergent abilities as model size increases, enhancing their capacity to understand context and handle long sequences (Xiong et al.,, 2023). This capacity growth allows LLMs to generate coherent and contextually accurate responses and enables various downstream applications, such as document summarization (Zhang et al.,, 2019; Zhang et al., 2024a, ), code generation (Chen et al., 2021b, ), and conversational AI (Bordes et al.,, 2016; OpenAI,, 2022), .
Despite their success in various applications, the generation of LLMs incurs significant expenses, which escalate with increasing model size and sequence length. Notably, both the training (Strubell et al.,, 2020; Hoffmann et al.,, 2022; Dong et al., 2024a, ) and inference (Ainslie et al.,, 2023) stages involve frequent generation by LLMs, further contributing to these costs. Consequently, efficient LLMs have gained popularity in recent years (Hu et al.,, 2021; Wan et al.,, 2023). To address these challenges, quantization (Frantar et al.,, 2022; Lin et al.,, 2024; Dettmers et al.,, 2024; Xu et al.,, 2023) and pruning methods (Frankle and Carbin,, 2018; Blalock et al.,, 2020) are employed to reduce model size. Additionally, managing long sequences presents another cost due to the transformer attention mechanism. The quadratic complexity of the attention mechanism results in substantial computational burden when dealing with long sequences, scaling poorly with sequence length. Therefore, effective management of long sequences is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario and aim to reduce the memory consumption associated with lengthy sequences.
Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size of each head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference. To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or performing quantization over the caches. It is well-acknowledged that token importance tends to be sparse. Consequently, KV eviction algorithms have been proposed to minimize the KV cache memory cost from dimension $S$ (Xiao et al., 2023b, ; Li et al.,, 2024; Zhang et al., 2024c, ; Leviathan et al.,, 2023). Additionally, inter-layer redundancy has been explored (Liu et al., 2024a, ; Wu and Tu,, 2024; Brandon et al.,, 2024) to address the layer dimension $L$. Despite these advances, existing methods have largely overlooked the channel dimension $D$. In this paper, we highlight that the magnitude across Key cache channel dimensions is significantly unbalanced. Additionally, we observe that the attention weights exhibit a low-rank structure. Based on these findings, we hypothesize that the channel dimension of the key cache exhibits redundancy. Consequently, we explore the redundancy of KV cache tensors specifically from dimension $D$, aiming to devise effective strategies that reduce costs without compromising performance.
In our paper, we propose a simple yet effective method – ThinK, for KV cache pruning. To pinpoint the least significant channels, we formulate the task as an optimization problem, aiming to minimize the loss in attention weights attributable to pruning. To effectively address this problem, we establish a novel query-dependent criterion that assesses the importance of each channel. Using this criterion, we then select the most critical channels in a greedy fashion. We evaluate ThinK using the LLaMA3 (Meta,, 2024) and Mistral (Jiang et al.,, 2023) models and validate its effectiveness across various long-sequence datasets. The results indicate that when paired with token eviction methods, ThinK not only achieves comparable or superior accuracy but also reduces KV cache memory costs by more than 20%.
Contributions.
This work pioneers the investigation into the sparsity structure of the channels in the KV cache. Specifically, we discovered that the activated key cache is sparse given a specific query. This insight allows us to prune the key cache channels using a query-induced norm. Inspired by this phenomenon, we introduce a plug-and-play method ThinK, the first channel pruning method specifically designed for KV cache. ThinK reduces the dimensionality of the cache channel, leading to linear savings in both memory and computational requirements. As plug-and-play techniques, our proposed ThinK is orthogonal to other KV cache compression schemes (e.g. KV cache eviction, quantization). Our extensive experiments demonstrate the remarkable efficiency of ThinK on LLaMA-3 and Mistral models. Furthermore, we explore the promising potential of extending ThinK to value cache pruning (ThinKV) in specific scenarios, highlighting the broad applicability of our method.
Figure 1: An illustration of the pruning procedure of ThinK. Within each head, scores are calculated for each channel, and only the top T channels out of D are selected for retention. A binary channel mask, along with the pruned keys, is subsequently stored in the cache memory.
2 Observations
We identify several key observations that motivate our approach to prune the channel of KV-Cache. Specifically, we visualize the magnitude of KV-cache and perform singular value decomposition of the attention in LLaMA3-8B model. We show the visualization for certain heads and layers for illustration purposes.
Magnitude of KV-Cache Channel
Figure 2 visualizes the absolute values of key and value cache across the tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis.. Consistent with previous findings (Lin et al.,, 2024; Xiao et al., 2023a, ; Liu et al., 2024b, ), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns. For instance, the absolute value around $50^{th}$ channel of key cache has much larger magnitudes than other channels for all tokens in layer $14$ (Figure 2 (a)). The same observation applies to $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 2 (c)). Given such an observation, Liu et al., 2024b proposed to perform quantization over the channels of the key cache based on such observations. In addition to channel quantization, the observation also suggests that it is feasible to prune away certain key cache channels that contribute less to the attention mechanism. Moreover, channel quantization and pruning techniques are orthogonal such that they can be utilized concurrently to enhance model efficiency.
(a) Key Cache ($14^{th}$ layer)
(b) Value Cache ($14^{th}$ layer)
(c) Key Cache ($20^{th}$ layer)
(d) Value Cache ($20^{th}$ layer)
Figure 2: Magnitude of key and value cache for LLaMA3-8B. The first head of layer 14 and layer 20 of LLaMA3-8B is selected to visualize the magnitude of the key and value caches. We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
Singular Value Analysis
We conducted singular value decomposition (SVD) (Demmel,, 1997) on the attention of the specified layers to investigate their principal components. The singular values derived from SVD capture the effective rank of the attention matrix, indicating how the information is distributed across different components.
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}$ $=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}$ (1)
Figure 3 (a) illustrates the energy distribution of the singular values, plotted in log scale to enhance the visibility of differences among them. Notably, only a few singular values exhibit high energy levels exceeding $0.01$ across all heads and layers, highlighting their relative significance. This observation aligns with previous findings (Bhojanapalli et al.,, 2021), where a small subset of singular values often captures most of the information in attention mechanisms. In addition, the rapid decay of the energy suggests that a low-rank approximation can effectively capture the essential information in the key cache.
Figure 3 (b), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy. These findings suggest that the attention matrix is inherently low-rank (Wang et al.,, 2020; Chen et al., 2021a, ; Dong et al., 2024b, ), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al.,, 2024).
(a) Energy of singular values
(b) Cumulative energy of singular values
Figure 3: The energy and cumulative energy of the singular values.
3 ThinK
Notations.
[…]
3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows:
$2\times B\times S\times L\times N\times D,$ (2)
where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cach size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al.,, 2017), such as LLaMA2-7B (Touvron et al., 2023b, ), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al.,, 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b, ; Zhang et al., 2024c, ; Li et al.,, 2024) and precision (Hooper et al.,, 2024; Liu et al., 2024b, ). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
Table 1: Performance comparison of pruning K cache by $l_{p}$ norm on LongBench.
Method $\lambda$ Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O 0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+$l_{1}$ 30 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
+$l_{1}$ 40 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
+$l_{2}$ 30 23.98 17.04 35.19 39.27 31.29 20.4 21.62 22.46 23.34 40.5 89.75 40.71 5.54 68.67 60.12 58.52 37.40
+$l_{2}$ 40 23.76 16.23 32.19 40.23 32.13 20.69 21.3 22.25 23.2 39.5 89.61 40.24 5.66 69.0 60.09 59.45 37.22
SnapKV 0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
+$l_{1}$ 30 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
+$l_{1}$ 40 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
+$l_{2}$ 30 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.5 90.33 40.31 5.7 68.42 62.65 60.07 40.28
+$l_{2}$ 40 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
Magnitude base Pruning: Based on the observations in Figure 2 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
$M_{n,d}=\big{\|}K[n,:,d]\big{\|}_{p}.$ (3)
Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head:
$I=\textbf{Top}_{T}(M,T)$ (4)
where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel. $n\in[1,N]$ and $d\in[1,D]$ are indicators of heads and channels in key cache. $I\in\mathbb{(Z^{+})}^{N\times T}$ stores the indicators of the top $T$ values in tensor $M$ per head.
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al.,, 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c, ) and SnapKV (Li et al.,, 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.
3.2 Query-Driven Pruning
For each head, the attention scores are computed using the queries and keys, and then applied to the values. The formula for the attention for head $i$ is:
$\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}$ (5)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. When pruning one channels of $\textbf{K}_{i}$, the corresponding channels in $\textbf{Q}_{i}$ will also be removed. We aim to find the optimal subset of channels to prune, denoted by the selection matrix $\textbf{S}\in\{0,1\}^{D\times D}$, where S is a diagonal matrix with binary entries ($1$ for keeping a channel, 00 for pruning it). To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights:
$\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}$ (6)
Given a pruning ratio $\lambda$, it can further expanded as:
$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$
subject to $\quad\text{trace}(\mathbf{S})=\lfloor(1-\lambda)D\rfloor$
$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$ (7)
For simplicity, we use greedy algorithm to optimize $\mathbf{S}$. To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores:
$\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F}$ (8)
$I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$ (9)
Here’s a detailed explanation of why it optimizes the selection matrix. The $\text{score}_{i}[j]$ measures the magnitude of the interaction between the query and key vectors for channel $j$ in each head $i$. By selecting channels with the highest interaction magnitudes, we aim to retain the most significant contributions to the attention mechanism. This criterion ensures that the selected channels preserve the primary information flow in the attention computation, thereby minimizing the loss of important information.
Observation Window. Based on the observations in SnapKV (Li et al.,, 2024) that the last window of input sequence recognizes highly similar attention allocation pattern with the actual generation. To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
Figure 4: Implementation during decoding.
3.3 Implementations
Following SnapKV, we focus on the long context input scenario. We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size. Additionally, we maintain a binary mask to indicate which channels have been pruned. Note that the memory overhead associated with this mask is negligible. Figure 4 illustrates one implementation of our method during the decoding stage. The pruned keys are first zero-filled using the mask to restore them to their original size, and then concatenated with the unpruned keys. This implementation can be effectively integrated with optimization techniques such as FlashAttention (Dao,, 2023). The second implementation involves initially pruning the Query using the mask. The pruned Query is then multiplied by the pruned Key, while the unpruned Query is applied to the unpruned Key. Subsequently, the two outputs are concatenated. In theory, this implementation has lower computational costs.
4 Experiment Results
In this section, we conduct extensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1. Next, we report the performance of our KV pruning method with two different KV compression methods in long-context scenarios in Section 4.2. Finally, we present the results of ablation experiments in Section 4.4.
4.1 Settings
Benchmark Datasets.
We perform evaluations of our method with state-of-the-art KV cache compression methods on two widely used benchmarks: LongBench and Needle-in-a-Haystack. LongBench (Bai et al.,, 2023) is designed to comprehensively evaluate LLM’s long context understanding capabilities. It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference. Needle-in-a-Haystack (Kamradt,, 2023) is a recent popular test challenge that requires models to accurately identify a small piece of information (“needle”) in a long document (“haystack”), where the needle is placed at a random position. This challenge can test if KV cache compression methods still retain the small piece of critical information.
Baseline Approaches.
The baseline methods include Heavy Hitter Oracle (H2O) and SnapKV, both of which are the state-of-the-art KV cache compression methods but use different strategies for selecting KV cache. H2O (Zhang et al., 2024c, ) is designed to reduce memory usage by dynamically balancing recent tokens and Heavy Hitter (H2) tokens, where H2 tokens are a small set of tokens that contribute most of the value when computing attention scores. SnapKV (Li et al.,, 2024) automatically compresses KV caches by selecting clustered important KV positions for each attention head. This is inspired by the fact that each attention head has a specific pattern, focusing on certain attention features during generation.
Table 2: Performance comparison of Key cache pruning on LLaMA3-8B by ThinK on LongBench.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
KV-size Full
ALL KV 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
KV-size 128
H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
+ThinK(40) 22.85 14.55 29.49 38.63 30.84 18.90 20.12 21.96 20.68 38.50 86.38 38.40 5.50 69.17 57.93 56.12 35.63
+ThinK(50) 23.47 14.06 28.67 38.35 30.21 17.87 19.69 21.94 19.95 38.50 87.14 38.07 4.92 69.50 57.99 56.66 35.44
[…]
+ThinK(40) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
+ThinK(50) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
KV-size 512
H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+ThinK(40) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
+ThinK(50) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
+ThinK(60) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
[…]
+ThinK(40) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
+ThinK(50) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
+ThinK(60) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
KV-size 1024
H2O 25.62 22.16 36.81 41.01 33.53 19.41 23.28 22.65 25.41 46.50 90.82 41.78 5.79 69.25 59.69 55.50 38.70
+ThinK(40) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
+ThinK(50) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
+ThinK(60) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
[…]
+ThinK(40) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
+ThinK(50) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
+ThinK(60) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
KV-size 2048
H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
+ThinK(40) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
+ThinK(50) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
+ThinK(60) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
[…]
+ThinK(40) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
+ThinK(50) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
+ThinK(60) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
Implementation Details.
In this paper, we use LLaMA3-8B-Instruct (Meta,, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al.,, 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al.,, 2020). Our ThinK aims to prune channels of the KV, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using one Nvidia A100. To fairly compare KV cache compression methods and KV cache compression integrated with ThinK, we use the same hyperparameters for both. For example, when comparing SnapKV and SnapKV integrated with ThinK, we set the maximum pooling kernel size to 7 and the observation window size to 32, using the same KV-size for both.
Table 3: Performance comparison of Key cache pruning on Mistral-7B by ThinK on LongBench.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
KV-size Full
ALL KV 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
KV-size 128
H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
+ThinK(40) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
+ThinK(50) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
+ThinK(60) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
[…]
+ThinK(40) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
+ThinK(50) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
+ThinK(60) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
KV-size 512
H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
+ThinK(40) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
+ThinK(50) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
+ThinK(60) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
[…]
+ThinK(40) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
+ThinK(50) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
+ThinK(60) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
KV-size 1024
H2O 23.67 28.55 46.40 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
+ThinK(40) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
+ThinK(50) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
+ThinK(60) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
[…]
+ThinK(40) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
+ThinK(50) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
+ThinK(60) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
KV-size 2048
H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.00 86.35 42.48 2.72 86.64 56.98 53.91 40.69
+ThinK(40) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
+ThinK(50) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
+ThinK(60) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
[…]
+ThinK(40) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
+ThinK(50) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
+ThinK(60) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52
4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and KV compression methods integrated with our channel pruning method of Key cache (ThinK) over two different base LLMs across various KV-sizes on LongBench. The following observations can be drawn: (1) Our method can further prune the channels of the key cache after compressing the KV cache with H2O and SnapKV. For the base model LLaMA3-8B, our method reduces memory usage and slightly improves performance for both H2O and SnapKV. For the base model Mistral-7B, our method reduces memory with only a slight drop in performance in some cases. (2) Comparing the results of SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm shown in Table 1, our proposed query-based channel pruning shows better performance when the pruning ratio is $40\%$ compared to no pruning. However, $l_{1}$ and $l_{2}$ norm methods cannot outperform the KV compression methods without pruning. (3) When the pruning ratio is small, performance tends to be better than when the pruning ratio is large. (4) When the KV-size is increased from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV cache size of $2048$ and a pruning ratio of $40\%$, our method can even outperform the LLaMA3-8B with a full KV cache. The above observations indicate that our method is agnostic to existing KV cache compression methods and can further improve their performance and memory reduction. Additionally, query-based pruning is more effective than $l_{1}$ and $l_{2}$ norm for channel pruning in LLMs.
4.3 Results on Needle-in-a-Haystack
Table 4 presents the results of the Needle-in-a-Haystack test based on the SnapKV (Li et al.,, 2024) approach with KV cache size of input prompt from $128$ to $2048$. With a modest pruning ratio $\lambda=40$%, the accuracy of ThinK is consistently better or comparable than the original SnapKV across both LLaMA3 and Mistral with different KV cache sizes. Such comparisons demonstrate that the proposed query-driven channel pruning can effectively retain the informative channels and discard the noisy ones. With a pruning ratio of $\lambda\geq 50$%, we observe an accuracy drop with a small KV-cache size, particularly for $128$ and $512$ across both LLaMA3 and Mistral. However, ThinK is still able to obtain comparable performance with SnapKV when we have a larger KV cache size (i.e., $1024$ and $2048$). Intuitively, a larger pruning ratio with a smaller KV cache size could tend to lose more important information compared to a larger KV cache size. In addition, the performance on larger KV cache size also suggests ThinK is robust for long context retrieval tasks.
Figure 6 (a)-(d) visualize the Needle-in-a-Haystack test accuracy against different token lengths and depth percent. The KV cache size is configured to $128$ and $1024$, with pruning ratio $\lambda$ equal to $40$% and $50$%, respectively. First, ThinK can maintain the original retrieval capability of SnapKV though the overall numerical accuracy is different (i.e., $77.8$ vs $78.6$ and $90.4$ vs $90.6$). Because ThinKachieves the same accuracy at most of the entries with different token limits and depths as Snap KV. In addition, ThinK also successfully retrieves some “needles” where SnapKV fails, leading to higher overall accuracy. The visualization results further demonstrate the robustness of ThinK in a fine-grained perspective and ThinK can enhance the original quantization approach.
Table 4: Needles-in-a-Haystack Test Results
Model Method $\bm{\lambda}$ KV-size
128 512 1024 2048
LLaMA3-8B SnapKV 00 79.6 90.2 91.2 91.7
SnapKV+ThinK 40 79.6 90.3 91.2 91.7
SnapKV+ThinK 50 77.4 89.6 91.0 91.7
Mistral-7B SnapKV 00 77.8 89.5 90.4 90.8
SnapKV+ThinK 40 78.6 90.1 90.6 90.9
SnapKV+ThinK 50 78.1 90.1 90.8 91.1
SnapKV+ThinK 60 75.9 89.2 90.6 91.1
4.4 Ablation Studies
Table 5: Performance comparison of pruning K cache using different recent-size.
Recent-Size Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O + ThinK($\lambda=40\%$)
0 24.40 27.50 45.42 35.17 24.45 13.02 27.65 23.88 26.86 53.50 86.06 41.73 3.01 83.42 55.12 51.32 38.91
32 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
128 25.69 30.93 48.32 39.63 26.08 16.82 27.18 23.92 26.62 53.50 86.39 42.96 3.29 86.39 56.77 53.60 40.51
SnapKV + ThinK($\lambda=40\%$)
0 24.94 28.58 45.78 39.59 25.40 15.92 29.50 24.05 26.72 70.00 85.60 41.38 2.97 84.00 55.27 52.39 40.76
32 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
128 25.75 32.49 48.61 41.01 27.18 19.14 28.79 24.64 26.77 70.00 86.37 42.77 3.61 87.13 57.19 54.39 42.24
Table 6: ablations with the same memory consumption.
Methods Memory(M) Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O
Vanilla 54.5 21.29 20.69 37.66 28.65 21.08 14.01 20.20 22.11 21.33 38.50 82.55 39.87 3.66 78.14 50.32 48.54 34.29
ThinK 54.4 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
Vanilla 208.0 22.13 23.83 43.24 30.92 23.36 14.56 22.92 22.77 24.23 41.50 85.04 41.26 3.02 86.03 54.91 50.50 36.89
ThinK 208.0 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
Vanilla 413.0 22.90 28.45 46.16 35.57 23.86 13.74 24.90 23.19 25.77 44.50 85.54 41.97 3.22 85.82 55.96 52.33 38.37
ThinK 412.8 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
Vanilla 822.5 25.51 30.23 48.23 39.72 25.56 16.75 26.98 23.81 26.47 50.50 86.43 42.09 2.78 85.57 57.4 53.42 40.09
ThinK 822.4 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
SnapKV
Vanilla 54.5 19.25 19.95 42.80 35.88 21.96 14.59 18.76 21.71 20.70 46.00 84.12 39.43 2.59 65.36 51.39 47.81 34.52
ThinK 54.4 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
Vanilla 208.0 23.31 27.45 48.85 38.77 23.93 16.50 23.44 23.63 24.13 66.00 86.05 41.00 2.62 87.01 56.13 52.60 40.09
ThinK 208.0 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.90 42.51 2.92 85.32 55.89 53.35 40.40
Vanilla 413.0 24.24 29.53 49.13 40.48 25.05 18.74 25.46 23.64 25.60 68.00 86.14 41.42 3.03 88.55 57.08 53.86 41.25
ThinK 412.8 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.30 56.71 54.19 41.62
Vanilla 822.5 24.84 31.90 48.16 41.32 26.77 19.49 28.23 24.63 26.41 70.00 86.32 41.83 2.91 88.06 56.98 53.74 41.97
ThinK 822.4 25.77 32.67 48.7 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
Impact of Different Recent Sizes.
Preserve the most recent KV embeddings (Zhang et al., 2024c, ; Li et al.,, 2024) is important for maintaining the performance of LLMs after KV cache compression. Note that a tradeoff exists: increasing the recent-size allows more infomation to be propagated, while increasing the cache size to be stored. To study its impacts, we evaluate the performance produced by three recent-size, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the basleline model. The results are summarized in Table 5. One can observe that a recent-size of 32 yields superior performance than 00 in terms of averaged score on LongBench which demonstrates the importance of keeping the most recent KVs. On the other hand, the performance of $32$ and $128$ is negligible, suggesting that maintaining the most recent 32 KVs suffices to preserve optimal performance.
Figure 5: Performance comparisons under the same KV cache memory consumption.
Comparison Under the Same Memory Consumption.
To ensure a fair comparison, we adjust the KV cache size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK. For example, the KV cache size of H2O with ThinK is $128$. Due to channel pruning of the key cache, the memory consumption of H2O with ThinK at a $128$ KV-size is lower than H2O at a 128 KV-size. Therefore, the KV-size of H2O is adjusted from $128$ to $109$. Table 6 and Figure 5 present the results of these comparisons on LongBench. As the results show, H2O or SnapKV with ThinK deliver superior performance compared to H2O or SnapKV without ThinK while maintaining the same memory usage. This demonstrates that KV cache compression methods, combined with our query-based channel pruning, can compress the KV cache more effectively and utilize memory more efficiently.
Pruning Channels of Both Key and Value Cache.
In our previous study, we evaluate methods with pruning channels of the key cache. In this part, we explore the impact of pruning channels of the value cache (Appendix A.1). Specifically, for KV cache compression methods, we prune channels of the key cache at one pruning ratio and channels of the value cache at another pruning ratio. Table 7 presents the results with LLaMA3-8B and Mistral-7B on LongBench. When testing on the base model LLaMA3-8B, H2O or SnapKV with key and value channel pruning can perform on par with H2O or SnapKV without channel pruning. In some cases, H2O or SnapKV with key and value channel pruning may even perform better. Pruning the channels of the value cache over the base model Mistral-7B experiences a slight performance drop. Despite this, note that KV cache with additional value channel pruning can further reduce memory usage.
Table 7: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench. H2O $+$ ThinK($\lambda_{1}$+$\lambda_{2}$) indicates that H2O has $\lambda_{1}$of the Key cache channels pruned and $\lambda_{2}$of the Value cache channels pruned.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
LLaMA3 8B KV-size 128
H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
+ThinK(30+30) 23.71 13.65 33.08 41.86 29.88 18.04 19.60 21.65 20.26 38.00 86.08 38.61 5.16 69.50 57.59 55.19 35.74
[…]
+ThinK(30+30) 21.86 13.79 33.26 40.93 29.39 19.22 18.81 21.30 19.26 41.50 87.00 37.95 5.78 69.50 57.84 55.62 35.81
KV-size 512
H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+ThinK(30+30) 22.83 17.57 34.18 42.67 33.52 19.95 21.17 22.23 22.82 38.50 90.11 39.08 5.21 69.0 59.99 56.83 37.23
[…]
+ThinK(30+30) 24.57 24.59 38.09 44.61 34.37 20.37 21.23 21.95 23.30 66.00 90.69 39.38 5.60 69.00 61.75 58.46 40.25
KV-size 2048
H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
+ThinK(30+30) 25.03 26.77 39.68 42.12 33.08 19.59 23.00 22.89 25.27 51.00 91.11 40.58 5.23 69.00 61.12 57.95 39.59
+ThinK(40+40) 24.87 24.31 37.77 43.13 34.42 19.60 21.67 22.70 24.52 49.00 90.81 39.28 6.00 69.00 61.81 58.08 39.19
[…]
+ThinK(30+30) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
+ThinK(40+40) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
Mistral 7B KV-size 128
H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
+ThinK(30+30) 20.71 21.49 38.01 30.66 22.28 13.87 20.13 22.45 21.07 38.50 82.20 38.69 2.94 78.56 51.55 48.28 34.46
[…]
+ThinK(30+30) 19.92 20.61 42.68 37.63 23.19 15.09 18.97 21.93 20.55 45.00 84.06 39.33 2.99 66.00 51.51 47.51 34.81
KV-size 512
H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
+ThinK(30+30) 22.36 24.26 44.77 30.47 22.94 14.96 22.63 22.90 23.73 41.50 85.30 40.21 3.08 80.07 54.48 50.96 36.54
+ThinK(30+10) 22.14 25.15 45.29 31.78 23.21 14.62 23.36 22.70 24.51 41.50 85.61 41.58 2.75 84.03 54.50 51.09 37.11
[…]
+ThinK(30+30) 24.10 27.04 47.76 38.66 25.45 17.51 22.64 22.81 23.91 66.00 86.62 39.91 3.36 82.24 55.96 52.81 39.80
+ThinK(30+10) 23.90 28.14 48.35 39.03 24.83 16.68 23.51 23.12 24.34 67.50 86.09 41.69 2.65 84.34 57.29 53.22 40.29
KV-size 1024
H2O 23.67 28.55 46.4 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
+ThinK(30+30) 23.65 26.54 47.00 35.52 24.79 17.15 23.64 23.12 25.20 44.00 86.38 41.67 3.46 80.14 56.53 52.86 38.23
+ThinK(30+10) 24.13 28.57 46.31 35.59 24.92 15.34 24.58 23.33 25.93 45.50 85.91 42.97 2.57 83.64 55.39 52.73 38.59
[…]
+ThinK(30+30) 25.29 29.25 49.17 41.25 25.75 19.37 24.64 23.02 25.27 69.00 86.70 40.92 3.29 82.06 57.15 54.15 41.02
+ThinK(30+10) 25.84 29.30 49.56 41.44 25.29 19.02 25.21 23.73 25.72 69.00 86.69 42.55 2.44 85.76 57.55 54.10 41.45
KV-size 2048
H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.0 86.35 42.48 2.72 86.64 56.98 53.91 40.69
+ThinK(30+30) 25.60 28.74 47.54 38.67 26.25 17.35 24.54 23.27 26.15 51.00 87.01 43.02 2.94 81.46 56.41 54.26 39.64
+ThinK(30+10) 25.64 30.65 48.95 40.42 26.43 16.65 26.76 23.51 26.59 52.50 86.53 43.45 2.66 83.96 56.55 53.83 40.32
[…]
+ThinK(30+30) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
+ThinK(30+10) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94
(a) KV-size 128, SnapKV, Acc.77.8
(b) KV-size 128, SnapKV+ThinK($40$) Acc.78.6
(c) KV-size 1024, SnapKV, Acc.90.4
(d) KV-size 1024, SnapKV+ThinK($50$), Acc.90.8
Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2.
5 Related Work
In scenarios involving long contexts, the most significant computational burden from the attention mechanism is the key-value (KV) cache. Reducing the KV cache is a high priority for optimizing deployment efficiency. System-level optimizations, such as FlashAttention (Dao,, 2023) and PagedAttention (Kwon et al.,, 2023), have been developed to address this issue. Additionally, algorithm-level optimizations are being explored to further enhance efficiency.
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b, ) maintains a few initial tokens and some recent tokens based on the observation of attention sink, which may result in the loss of important information carried by the dropped tokens. H2O (Zhang et al., 2024c, ) retains only a small portion of the tokens by greedily dropping tokens based on their contributions to the cumulative attention. SnapKV (Li et al.,, 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al.,, 2023) adaptively evicts tokens from attention heads that emphasize local contexts. This approach focuses on discarding non-special tokens centered on special tokens, while the standard KV cache is used only for attention heads that broadly attend to all tokens. PyramidKV (Zhang et al., 2024b, ) and PyramidInfer (Yang et al.,, 2024) considers adjusting the KV cache size across different layers by allocating more cache in the lower layers and less in the higher ones.
KV Cache Quantization. SmoothQuant (Xiao et al., 2023a, ) can quantize the KV cache to 8-bit with minimal performance degradation. Q-Hitter (Zhang et al., 2024c, ) uses accumulated attention scores and ‘Quantization Friendliness’ metrics to identify tokens that are essential for maintaining the generalization capabilities of LLMs and are suitable for KV cache quantization. Some studies have found that the key cache and value cache should be quantized differently (Liu et al., 2024b, ; Hooper et al.,, 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
Structured Pruning of LLMs. Structured pruning  (Ma et al.,, 2023; Ding et al.,, 2023) of LLMs traditionally focuses on removing unimportant layers, heads, and hidden dimensions, which often results in significant performance degradation. In contrast, our methodology preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels based on data dependant criterion, our approach greatly reduce the key cache-size with negligible performance loss.
6 End Note and Future Direction
Motivated by the observation that certain channels have significantly larger magnitudes compared to others, and the singular value analysis indicates that the key cache is inherently low-rank (Singhania et al.,, 2024), we propose ThinK to perform pruning over the key cache channel. The proposed pruning strategy is query-dependant and optimized based on the attention scores, ensuring that essential information is retained for each input query. In addition, ThinK can be seamlessly integrated with other popular token-level KV cache quantization techniques (Li et al.,, 2024; Zhang et al., 2024c, ), further enhancing inference efficiency. Extensive experiments on LongBench (Bai et al.,, 2023) and Needle-in-a-Haystack tests with two foundation models demonstrate the effectiveness and robustness of our query-dependent channel pruning method. Our approach achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%. Our analysis indicates that ThinK can maintain superior performance over baselines with a smaller KV cache size under equivalent memory consumption conditions.
Future work will focus on strategies to increase the pruning ratio without compromising model performance, and exploring pruning techniques for value caches. We also plan to evaluate channel importance from a more sophisticated and compositional perspective, incorporating both token-level and channel-level information to calculate importance.
Appendix A Additional Experimental Details
A.1 Value Cache Pruning
Similar to the approach used for the Key cache, the pruning of channels in the Value cache can be guided by two primary criteria: magnitude-based pruning and query-driven pruning. We find that query-driven pruning is still better than magnitude based pruning.
$\text{Score}_{v,i}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})[j]=\|\text{softmax}(\frac{\textbf{Q}_{i}[-S_{\text{obs}}:]\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}[:,j]\|_{F}$ (10)
$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (11)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache. Then, only top T channels are retained.